OpenStudy (elisuzsmith):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (zale101):

use point slope formula

OpenStudy (anonymous):

equation of a line through a point (x1,y1) and slope m is y-y1=m(x-x1)

OpenStudy (zale101):

\[y-y_1=m(x-x_1)\]

OpenStudy (zale101):

where m is the slope and the (x-x1) is the given point

OpenStudy (zale101):

the Slope (m)=3 and x1=8 and y1=-8 so plug in into that formula i provided above

OpenStudy (elisuzsmith):

y – 8 = 3 (x + 8)

OpenStudy (zale101):

no y-(-8)=3(y+8) rewrite y-(-8) to y+8

OpenStudy (zale101):

remember that negative and negative is a positive answer

OpenStudy (zale101):

it should be written like this y+8=3(y-8)

OpenStudy (zale101):

The point–slope form of the equation for a line passing through the point (x1,y1) is given by: \[y-y_1=m(x-x_1)\] Assuming (8, –8) = (x1, y1) and m = 3 and substituting the values, we get : \[y-(-8)=3(x-8)\] so simplify : \[y+8=3x-8\]